Germaine paid $599.20 for new tires with 7% sales tax included. What was the price of the tires? Please show all work.

Found 2 solutions by stanbon, Rcmathtutor:
Answer by stanbon(75887)

Germaine paid $599.20 for new tires with 7% sales tax included.
What was the price of the tires?
Equation:
x - 0.07x = 599.20
0.93x = 599.20
x = $644.30
--------------
Cheers,
Stan H.

Answer by Rcmathtutor(12)

Let X be the original, pre-tax price.
Pre-tax price + sales tax = price with tax included
X + 0.07X = 599.20
1.07X = 599.20
Divide both sides by 1.07
X = 599.20 / 1.07
X = 560
The original, pre-tax price, was $560
Proof:
560 + (560)(0.07) = 599.20